# Exciton-Polariton Dynamics in Multilayered Materials

### v1

@@ Abstract
Results and Discussion. Here we model the exciton-phonon part of the system based on a multilayered perovskite material [22], which has on-site energy $\epsilon_{0}=3.2$ eV, a phonon frequency of $\omega=1440$ cm-1 and a hopping parameter of $\tau=400$ cm-1 with an interlayer spacing of $4$ nm and a lattice spacing of $a=1.2$ nm. However, the conclusions of this work are generic and is applicable to a wide variety of organic and inorganic exciton-polaritons. Further details of the parameters used in this work are provided in the SI.
@@
Fig. 2a-b present the angle-resolved polariton spectra computed using our quantum dynamical approach (see details in the SI) for a single-layer material coupled to cavity radiation modes. The theoretical polariton dispersion curves presented in Fig. 1c, which are obtained by diagonalizing $\hat{H}_{\mathrm{EP}}$ (see Eq. 12), feature smooth polariton curves, as expected. In contrast, phonon interaction with the exciton-polaritons introduces vibronic structure to the polariton dispersion, as can be seen in Fig. 2a. Increase in the phonon coupling $\gamma=2\gamma_{0}$ (where $\gamma_{0}=1.1\sqrt{\omega^{3}}$), in Fig. 2b makes the vibronic structure more prominent. The existence of these vibronic bands and their relation to phonon coupling $\gamma$ can be, very crudely, explained due to the Franck–Condon overlaps between displaced harmonic oscillator states of the phonons. In our recent work [8], we have presented a microscopic theory, where these successive vibronic bands describe excitations to a phonon field, which provides a quantitatively accurate description.
Figure 3: Comparing exciton-polariton dynamics in single-layer vs. multilayered materials ($M=25$) with the initial excitation at $E_{0}=2.64$ eV. (a) Time-dependent exciton-polariton density for a single-layer material with no phonon couplings, i.e. $\gamma=0$. (b)-(c) Same as (a) but in the presence of phonon couplings $\gamma=2\gamma_{0}$ in (b) a single and (c) multilayered material. (d) Site-resolved exciton density at $t=900$ fs and $\gamma=2\gamma_{0}$. To determine the multilayer exciton populations in (d), we summed the density contributions from all layers at each site. (e)-(h) Site and layer-resolved exciton density for (e)-(f) a single layer and (g)-(h) multilayered material at $t=0$ and $t=900$ fs with the same parameters as in (d). ($\gamma_{0}\approx 5.8\times 10^{-4}$ a.u.)
@@
Importantly, the vibronic structure in polariton dispersion directly shows up in the group velocities obtained via direct quantum dynamics simulations. In our quantum dynamics simulations, we model a resonant excitation in the lower polariton at the energy $E_{0}$ by selecting an exciton-polariton wavefunction $|\Psi(0)\rangle=\sum_{k\in\mathcal{E}_{\mathrm{exc}}}\tilde{c}_{k}\hat{P}_{k,-}^{\dagger}|\bar{0}\rangle$, which is energetically localized within an energy window $\mathcal{E}_{\mathrm{exc}}$ (set to be $50$ meV around $E_{0}$). We choose the coefficients $\tilde{c}_{k}$ via a Monte Carlo approach (see SI) to ensure that $|\Psi(0)\rangle$ is also spatially localized at the center in the plane of the materials (in the $\vec{x}$ direction).
@@
Fig. 2c-d presents the group velocities extracted from the time-dependent excitonic density $\langle\Psi(t)|\hat{X}_{n,m}^{\dagger}\hat{X}_{n,m}|\Psi(t)\rangle$ at short times ($t<300$ fs) with different numbers of layers of the material coupled to the cavity radiation. In Fig. 2c we keep each layer’s coupling to the radiation $\Omega_{k}$ a constant. As a result, the overall Rabi-splitting $\sqrt{\mathcal{S}}\Omega_{k}$ roughly scales with $\sim\sqrt{M}$ where $M$ is the number of layers. The group velocities obtained numerically (solid lines with filled circles) show an oscillatory pattern due to the presence of the vibronic structure in the exciton-polariton dispersion. An increase in the number of layers, from $M=1$ (yellow) to $M=5$ (light red), and $M=10$ (dark red), leads to an overall redshift due to the increase in the Rabi-splitting. Notably, the computed group velocities are much smaller (renormalized) than the theoretical group velocity $\frac{d}{dk}\omega_{-,k}$ (i.e. slope of the dispersion) in bare exciton-polariton. This renormalization effect has been observed in multiple recent experiments [2, 3, 6]. To clearly analyze the role of multiple layers in the exciton-polariton dynamics, in Fig. 2d and the rest of the results presented in this work we keep $\sqrt{\mathcal{S}}\Omega_{k}$ a constant.
@@
Fig. 2d presents the exciton-polariton group velocities in a multilayered material (with $M=25$) compared to a single-layered material at various phonon couplings. Fig. 2d reveals that for the same effective Rabi splitting, $\sqrt{\mathcal{S}}\Omega_{k}$, the exciton-polariton group velocity in a multilayered material could be substantially higher than the single-layered material. This is further illustrated in Fig. 2e which presents the enhancement of group velocity as a function of the number of layers placed inside an optical cavity. We find that the exciton-polariton density in multilayered material can propagate with velocities upwards of $\sim$50higher than in a single-layer material.
@@
This is surprising because the number of phonon modes interacting with the exciton-polariton subsystem is significantly higher in a multilayered material compared to a single-layer material. However, despite the increase in the number of phonon modes, which one expects to provide additional sources of dissipation and fluctuation, the exciton-polariton propagates faster with a longer decoherence time. We find that this effect originates from a synchronization of phonon fluctuations in the collective bright layer leading to the suppression of phonon-induced decoherence and enhanced group velocity.
@@
Phonon fluctuation synchronization effect. To provide a microscopic understanding of this effect, we analyze the term $\hat{H}_{\mathrm{ex-phn}}$ in the light-matter Hamiltonian. We introduce the bright exciton operators in real space, defined as $\hat{X}^{\dagger}_{n,b}=\frac{1}{\sqrt{\mathcal{S}}}\sum_{m}\sin(k_{y}\cdot y_{m})\hat{X}^{\dagger}_{n,m}=\frac{1}{\sqrt{N}}\sum_{n}\hat{X}^{\dagger}_{k,b}e^{-ik\cdot x_{n}}$. Since only the bright exciton operators $\{\hat{X}^{\dagger}_{n,b}\}\equiv\{\hat{X}^{\dagger}_{k,b}\}$ directly contribute (see Eq. 12) in the dynamics of the exciton-polaritons, we focus on the exciton-phonon coupling term in the subspace defined by the projection operator $\mathcal{P}=\sum_{n}\hat{X}^{\dagger}_{n,b}\hat{X}_{n,b}=\sum_{n}|n,b\rangle\langle n,b|$, which is written as
$\mathcal{P}\left(\hat{H}_{\mathrm{ex-phn}}\right)\mathcal{P}$ $=\gamma\sum_{n}\hat{X}^{\dagger}_{n,b}\hat{X}_{n,b}\left(\frac{1}{{\mathcal{S}}}\sum_{m}{q}_{n,m}\sin^{2}(k\cdot y_{m})\right)$
@@
where $\tilde{q}_{n}=\frac{1}{{\mathcal{S}}}\sum_{m}{q}_{n,m}\sin^{2}(k\cdot y_{m})$. Therefore, exciton-polariton dynamics effectively interact with the collective phonons $\tilde{q}_{n}$ that are delocalized over all the layers. Importantly, the $\tilde{q}_{n}$ distribution is much narrower than the distribution of ${q}_{n,m}$ effectively exhibiting a phonon fluctuation synchronization in multilayered materials. This phonon fluctuation synchronization can also be viewed as an effective reduction in temperature (shrinking the thermal distribution of the phonons) leading to enhanced coherent transport of exciton-polaritons.
@@
Fig.2f presents the distribution of the excitonic on-site energy compared to the multilayered material. Due to the fluctuation synchronization discussed above, the fluctuations in the effective excitonic on-site energy $\tilde{\epsilon}_{n}={\epsilon}_{0}+\gamma\tilde{q}_{n}$ in the collective bright layer is much less compared to the on-site energies ${\epsilon}_{n}={\epsilon}_{0}+\gamma{q}_{n}$ in the single-layer scenario. As a result, the exciton-polariton propagation is enhanced as hopping between neighboring sites is easier, which is schematically illustrated in Fig.2g-h.
@@
To test our phonon fluctuation synchronization hypothesis, in Fig. 2i, we perform quantum dynamical simulations in multilayered materials by sampling the phonons in a constrained fashion, i.e. ${q}_{n,m}(t=0)={q}_{n,1}(t=0)$, such that the initial position of a phonon at a site $n$ is the same regardless of the layer. As a result, $\tilde{q}_{n}={q}_{n,1}$, which means that the fluctuations experienced by the collective bright layer in a multilayered material is the same as in the single-layer material. The group velocities presented in Fig. 2i show that the dynamics of the multilayered material under constrained sampling are identical to the single-layer material, corroborating our phonon fluctuation synchronization hypothesis.
@@
Fig. 3 presents time-dependent exciton-polariton density (exciton population at each site $n$) to illustrate how the phonon fluctuation synchronization in multilayered materials can enable coherent transport. Fig. 3a presents the time-dependent exciton population in a single-layer material in the absence of phonon ($\gamma=0$). As expected, the exciton-polariton population, initially prepared at the center, propagates ballistically. In the presence of strong phonon interaction ($\gamma\neq 0$) the exciton-polariton propagation in a single-layer material, presented in Fig. 3b, appears incoherent, exhibiting diffusive transport. In contrast, the exciton-polariton transport in multilayered material at the same exciton-phonon coupling $\gamma$, presented in Fig. 3c, shows ballistic coherent propagation of the exciton-polariton wavefront. This remarkable effect can be clearly seen in Fig. 3d which presents the exciton-polariton density at $t=900$ fs. We see that the wavefront in multilayered material propagates coherently, which is missing in the single-layer case.
@@
Fig. 3g-h further illustrates that the exciton-polariton wavefront is delocalized over all the layers, with Fig. 3e-f presenting the corresponding results in the single-layer scenario. This shows how delocalization of exciton-polariton density in the transverse direction exhibits enhanced transport in the longitudinal direction (or in the plane of the material). We find that this phonon fluctuation synchronization effect is quite ubiquitous, showing up in a wide range of parameter regimes. We present more exciton density propagation data for different values of $E_{0}$ and $\gamma$ in the SI.
Figure 4: Effect of cavity photon loss on mean-square displacement (MSD) and purity in multilayered materials. MSD in single-layer material vs. multilayered material ($M=25$) (a) in the absence ($t_{c}=\infty$) and (b) presence ($t_{c}=263$ fs) of cavity photon loss. Purity in the absence of photon loss for different number of layers (M) and the initial excitation at (c) $E_{0}=2.58$ eV and (d) $E_{0}=2.78$ eV. Total purity (Tr$[\hat{\rho}_{\text{T}}^{2}(t)]$) along with exciton-polariton (Tr$[\hat{\rho}_{\text{EP}}^{2}(t)]$) and ground state (Tr$[\hat{\rho}_{\text{g}}^{2}(t)]$) contributions to total purity in the presence of cavity photon loss ($t_{c}=263$ fs) and initial excitation at $E_{0}=2.58$ eV for (e) single layer and (f) 15 layers. The ratio of the purity of multilayered materials ($M>1$) relative to the purity of single-layer material ($M=1$) with no cavity loss and initial excitation at (g) $E_{0}=2.58~{}\text{eV}$ and (h) $E_{0}=2.78~{}\text{eV}$. The ratio of exciton-polariton purity of 15-layer material relative to the single-layer material for different cavity photon lifetimes and initial excitation at (i) $E_{0}=2.58~{}\text{eV}$ and (j) $E_{0}=2.78~{}\text{eV}$. ($\gamma_{0}\approx 5.8\times 10^{-4}$ a.u.)
@@
Fig. 4a-b presents the excitonic mean square displacement (MSD) for an initially prepared exciton-polariton at an energy window of $2.64\pm 0.025$ eV. The excitonic MSD is computed as [2, 9, 29]
@@
where $P_{n}(t)=\left\langle\frac{1}{{\mathcal{N}}}\langle\Psi(t)|\hat{X}^{\dagger}_{n}\hat{X}_{n}|\Psi(t)\rangle\right\rangle_{\mathrm{MFE}}$ is excitonic density at the site $n$ with $\mathcal{N}=\sum_{n}\langle\Psi(t)|\hat{X}^{\dagger}_{n}\hat{X}_{n}|\Psi(t)\rangle$ a normalization constant. In the absence of phonons ($\gamma=0$), MSD is quadratic (dark blue line in Fig. 4a) in time, i.e. ballistic, as expected. In the presence of phonon interactions, ($\gamma\neq 0$) MSD is quadratic till $\sim 300$ fs and becomes almost linear beyond this time, thereby exhibiting diffusive transport as phonon-induced decoherence sets in. For the same phonon coupling, the MSD appears to be ballistic for a longer time in the case of multilayered material, suggestive of extended coherence lifetime due to the phonon fluctuation synchronization discussed above. Clearly the MSD reaches a higher value in the case of a multilayered material compared to a single-layer material.
@@
Fig. 4b presents MSD in the presence of cavity loss that accounts for the finite lifetime of a photon trapped inside an imperfect optical cavity. Overall, the trends are the same as in Fig. 4a but with an overall lower MSD. These results agree with recent theoretical and experimental works which demonstrate that MSD is suppressed in the presence of cavity loss [29, 9, 6].
@@
While ballistic transport is suggestive of coherent dynamics, it only provides indirect evidence of decoherence suppression. To provide a quantitative measure and a direct confirmation of decoherence suppression, we compute the purity [42, 43, 44] which is defined as $\mathrm{Tr}[\hat{\rho}_{\mathrm{EP}}^{2}(t)]$ where $\hat{\rho}_{\mathrm{EP}}(t)$ is the reduced density matrix describing the exciton-polariton subsystem. Within our mixed-quantum classical simulations, the exciton-polariton reduced density matrix is computed as $\big{\langle}|\Psi(t)\rangle\langle\Psi(t)|\big{\rangle}_{\mathrm{MFE}}$. In the SI we provide an efficient approach to computing the purity in our mixed quantum-classical simulations.
@@
Fig. 4c-d presents the time-dependent purity $\mathrm{Tr}[\hat{\rho}_{\mathrm{EP}}^{2}(t)]$ at two different windows of initial excitation $E_{0}=2.58$ and $2.78$ eV. In the single-layer scenario (violet solid line), at times $t=0$ the purity is one, which indicates a pure (coherent) state. At longer times, $\sim t=600$ fs the purity decays toward zero, indicating phonon-induced decoherence of the exciton-polariton subsystem. Fig. 4c-d, shows that an increase in the number of layers significantly suppresses the decay of the purity, directly depicting an extended coherence lifetime in multilayered materials. Overall, we find that multilayered materials can enhance purity by up to a factor of $\sim 8$ as shown in Fig. 4g-h which presents the ratio of purity in a multilayered material to a single-layer material.
@@
Fig. 4e-f presents the time-dependent purity in the presence of cavity loss. Note that in the presence of cavity loss, the population in the exciton-photon subspace $\{|1_{k}\rangle,|n,m\rangle\}$ leaks out to the ground state $|\bar{0}\rangle$. As a result, the total reduced density matrix (when tracing over the phonon degree of freedom) can be defined as $\hat{\rho}_{\mathrm{T}}(t)=\hat{\rho}_{\mathrm{EP}}(t)+\hat{\rho}_{\mathrm{g}}(t)=\sum_{\alpha,\beta}\rho_{\alpha\beta}(t)|\alpha\rangle\langle\beta|+\rho_{\mathrm{gg}}(t)|\bar{0}\rangle\langle\bar{0}|$ where $\{|\alpha\rangle,|\beta\rangle\}\in\{|1_{k}\rangle,|n,m\rangle\}$. As a result, the total purity can be written down as
$\mathrm{Tr}[\hat{\rho}_{\mathrm{T}}^{2}(t)]=\mathrm{Tr}[\hat{\rho}_{\mathrm{EP}}^{2}(t)]+\mathrm{Tr}[\hat{\rho}_{\mathrm{g}}^{2}(t)].$ (18)

### v2

@@ Abstract
Results and Discussion. Here we present results for exciton-polariton transport where the exciton-phonon part of the system is modeled based on a multilayered perovskite material [23], which has on-site energy $\epsilon_{0}=3.2$ eV, a phonon frequency of $\omega=1440$ cm-1 and a hopping parameter of $\tau=400$ cm-1 with an interlayer spacing of $4$ nm and a lattice spacing of $a=1.2$ nm. However, the conclusions of this work are generic and apply to a wide variety of organic and inorganic exciton-polaritons. Further details of the parameters used in this work are provided in the SI.
@@
Fig. 2a-b present the angle-resolved polariton spectra computed using our quantum dynamical approach (see details in the SI) for a single-layer material coupled to cavity radiation modes. The theoretical polariton dispersion curves presented in Fig. 1c, which are obtained by diagonalizing $\hat{H}_{\mathrm{EP}}$ (see Eq.  Exciton-Polariton Dynamics in Multilayered Materials), feature smooth polariton curves, as expected. In contrast, phonon interaction with the exciton-polaritons introduces vibronic structure to the polariton dispersion, as can be seen in Fig. 2a. Increase in the phonon coupling $\gamma=2\gamma_{0}$ (where $\gamma_{0}=1.1\sqrt{\omega^{3}}$), in Fig. 2b makes the vibronic structure more prominent. The existence of these vibronic bands and their relation to phonon coupling $\gamma$ can be, very crudely, explained due to the Franck–Condon overlaps between displaced harmonic oscillator states of the phonons. In our recent work [8], we have presented a microscopic theory, where these successive vibronic bands describe excitations to a phonon field, which provides a quantitatively accurate description.
@@
Importantly, the vibronic structure in polariton dispersion is directly reflected in the group velocities obtained via direct quantum dynamics simulations. In our quantum dynamics simulations, we model a resonant excitation in the lower polariton at the energy $E_{0}$ by selecting an exciton-polariton wavefunction $|\Psi(0)\rangle=\sum_{k\in\mathcal{E}_{\mathrm{exc}}}\tilde{c}_{k}\hat{P}_{k,-}^{\dagger}|\bar{0}\rangle$, which is energetically localized within an energy window $\mathcal{E}_{\mathrm{exc}}$ (set to be $50$ meV around $E_{0}$). We choose the coefficients $\tilde{c}_{k}$ via a Monte Carlo approach (see SI) to ensure that $|\Psi(0)\rangle$ is also spatially localized at the center in the plane of the materials (in the $\vec{e}_{x}$ direction).
@@
Fig. 2c-d presents the group velocities extracted from the time-dependent excitonic density $\langle\Psi(t)|\hat{X}_{n,m}^{\dagger}\hat{X}_{n,m}|\Psi(t)\rangle$ at short times ($t<0.3$ ps) with different numbers of layers of the material coupled to the cavity radiation. In Fig. 2c we keep each layer’s coupling to the radiation $\Omega_{k}$ a constant. As a result, the overall Rabi-splitting $\sqrt{\mathcal{S}}\Omega_{k}$ roughly scales with $\approx\sqrt{M}$ where $M$ is the number of layers. The group velocities obtained numerically (solid lines with filled circles) show an oscillatory pattern due to the presence of the vibronic structure in the exciton-polariton dispersion. An increase in the number of layers, from $M=1$ (yellow) to $M=5$ (light red), and $M=10$ (dark red), leads to an overall redshift due to the increase in the Rabi-splitting. Notably, the computed group velocities are much smaller (renormalized) than the theoretical group velocity $\frac{d}{dk}\omega_{-,k}$ (i.e. slope of the dispersion) in bare exciton-polariton. This renormalization effect has been observed in multiple recent experiments [2, 3, 6]. To clearly analyze the role of multiple layers in the exciton-polariton dynamics, in Fig. 2d and the rest of the results presented in this work we keep $\sqrt{\mathcal{S}}\Omega_{k}$ a constant.
@@
Fig. 2d presents the exciton-polariton group velocities in a multilayered material (with $M=25$) compared to a single-layered material at various phonon couplings. Fig. 2d reveals that for the same effective Rabi splitting, $\sqrt{\mathcal{S}}\Omega_{k}$, the exciton-polariton group velocity in a multilayered material could be substantially higher than the single-layered material. This is further illustrated in Fig. 2e which presents the enhancement of group velocity as a function of the number of layers placed inside an optical cavity. We find that the exciton-polariton density in multilayered material can propagate with velocities upwards of $\approx$25% higher than in a single-layer material.
@@
This is surprising because the number of phonon modes interacting with the exciton-polariton subsystem is significantly higher in a multilayered material compared to a single-layer material. However, despite the increase in the number of phonon modes, which one expects to provide additional sources of dissipation and fluctuation, the exciton-polariton propagates faster with a longer decoherence time. We find that this effect originates from a synchronization of phonon fluctuations in the collective bright layer, leading to the suppression of phonon-induced decoherence and enhanced group velocity.
@@
Phonon fluctuation synchronization effect. To provide a microscopic understanding of phonon fluctuation synchronization effect, we analyze the term $\hat{H}_{\mathrm{ex-phn}}$ in the light-matter Hamiltonian. We introduced the bright exciton operators in real space as $\hat{X}^{\dagger}_{n,b}=\frac{1}{\sqrt{\mathcal{S}}}\sum_{m}\sin(k_{z}z_{m})\hat{X}^{\dagger}_{n,m}=\frac{1}{\sqrt{N}}\sum_{n}\hat{X}^{\dagger}_{k,b}e^{-ikx_{n}}$. Since only the bright exciton operators $\{\hat{X}^{\dagger}_{n,b}\}\equiv\{\hat{X}^{\dagger}_{k,b}\}$ directly contribute (see Eq.  Exciton-Polariton Dynamics in Multilayered Materials) in the dynamics of the exciton-polaritons, we focus on the exciton-phonon coupling term in the subspace defined by the projection operator $\mathcal{P}=\sum_{n}\hat{X}^{\dagger}_{n,b}\hat{X}_{n,b}=\sum_{n}|n,b\rangle\langle n,b|$, which is written as
$\mathcal{P}\left(\hat{H}_{\mathrm{ex-phn}}\right)\mathcal{P}$ $=\gamma\sum_{n}\hat{X}^{\dagger}_{n,b}\hat{X}_{n,b}\left(\frac{1}{{\mathcal{S}}}\sum_{m}{q}_{n,m}\sin^{2}(k_{z}z_{m})\right)$
@@
where $\tilde{q}_{n}=\frac{1}{{\mathcal{S}}}\sum_{m}{q}_{n,m}\sin^{2}(k_{z}z_{m})$. Note that we do not make such assumption when performing any numerical calculations. Therefore, exciton-polariton dynamics effectively interact with the collective phonon fluctuations $\tilde{q}_{n}$ that are delocalized over all the layers. Importantly, the $\tilde{q}_{n}$ distribution is much narrower than the distribution of ${q}_{n,m}$, effectively exhibiting a phonon fluctuation synchronization in multilayered materials. This phonon fluctuation synchronization effect can also be viewed as an effective reduction in temperature (shrinking the thermal distribution of the phonons), leading to enhanced coherent transport of exciton-polaritons (see the SI for more details).
@@
Fig.2f presents the distribution of the excitonic on-site energy in single layer materials compared to the multilayered materials. Due to the phonon fluctuation synchronization effect discussed above, the fluctuations in the effective excitonic on-site energy $\tilde{\epsilon}_{n}={\epsilon}_{0}+\gamma\tilde{q}_{n}$ in the collective bright layer of multilayer materials is much less compared to the on-site energies ${\epsilon}_{n}={\epsilon}_{0}+\gamma{q}_{n}$ in the single-layer scenario. As a result, the exciton-polariton propagation is enhanced as the exciton hopping between neighboring sites is easier, which is schematically illustrated in Fig.2g-h.
@@
To test our phonon fluctuation synchronization hypothesis, in Fig. 2i, we performed mixed-quantum-classical simulations in multilayered materials by sampling the phonons in a constrained fashion, i.e. ${q}_{n,m}(t=0)={q}_{n,1}(t=0)$, such that the initial position of a phonon at a site $n$ is the same regardless of the layer index. As a result, $\tilde{q}_{n}={q}_{n,1}$, which means that the fluctuations experienced by the collective bright layer in a multilayered material are the same as in the single-layer material. The group velocities presented in Fig. 2i show that the dynamics of the multilayered material under constrained sampling are identical to the single-layer material, corroborating our phonon fluctuation synchronization hypothesis.
@@
Fig. 3 presents time-dependent exciton-polariton density (exciton population at each site $n$) to illustrate how the phonon fluctuation synchronization in multilayered materials can enable coherent transport. Fig. 3a presents the time-dependent exciton-polariton population in the absence of phonon ($\gamma=0$) for both single-layer and multilayer materials. As expected, the exciton-polariton population, initially prepared at the center, propagates ballistically. In the presence of strong phonon interaction ($\gamma\neq 0$) the exciton-polariton propagation in a single-layer material, presented in Fig. 3b, appears incoherent, exhibiting diffusive transport. In contrast, the exciton-polariton transport in multilayered material at the same exciton-phonon coupling $\gamma$, presented in Fig. 3c, shows ballistic coherent propagation of the exciton-polariton wavefront. This remarkable effect can be clearly seen in Fig. 3d, which presents the exciton-polariton density at $t=0.75$ ps. We see that the wavefront in a multilayered material propagates coherently, which is missing in the single-layer case.
@@
Fig. 3g-h further illustrates that the exciton-polariton wavefront is delocalized over all the layers, with Fig. 3e-f presenting the corresponding results in the single-layer scenario. This shows how delocalization of exciton-polariton density in the transverse direction exhibits enhanced transport in the longitudinal direction (or in the plane of the material). We find that this phonon fluctuation synchronization effect is quite ubiquitous, showing up in a wide range of parameter regimes. We present more exciton-polariton density propagation data for different values of $E_{0}$, $\Omega_{0}$, and $\gamma$ in the SI.
[…]
Figure 5: Comparing exciton-polariton dynamics in single-layer vs. multilayered materials ($M=10$) with the initial excitation at $E_{0}=2.43$ eV in a three-dimensional cavity. (a)-(c) Time-dependent exciton-polariton population (density) in the absence of exciton-phonon coupling for both single-layer and multilayer materials at three different times, (d)-(f) same as (a)-(c) but for the single-layer material and in the presence of exciton-phonon coupling ($\gamma=2\gamma_{0}$), (g)-(i) same as (a)-(c) but for the multilayered material ($M=10$) and in the presence of exciton-phonon coupling ($\gamma=2\gamma_{0}$). To determine the multilayer exciton-polariton populations, we summed the density contributions from all layers at each site. Population for longer times for all cases can be found in SI. ($\gamma_{0}\approx 5.8\times 10^{-4}$ a.u., $\Omega_{0}=330$ meV, and $N_{x}=N_{y}=1000$.)
@@
Fig. 4a-b presents the excitonic mean square displacement (MSD) of the exciton-polariton initially prepared at an energy window of $2.62\pm 0.025$ eV. The excitonic MSD is computed as [2, 9, 30]
@@
where $P_{n}(t)=\left\langle\frac{1}{{\mathcal{N}}}\langle\Psi(t)|\hat{X}^{\dagger}_{n}\hat{X}_{n}|\Psi(t)\rangle\right\rangle_{\mathrm{MFE}}$ is excitonic density at the site $n$ with $\mathcal{N}=\sum_{n}\langle\Psi(t)|\hat{X}^{\dagger}_{n}\hat{X}_{n}|\Psi(t)\rangle$ a normalization constant. In the absence of phonons ($\gamma=0$), MSD is quadratic in time (dark blue line in Fig. 4a), i.e. ballistic, as expected. In the presence of exciton-phonon coupling ($\gamma\neq 0$), MSD is quadratic in time till $\approx 0.3$ ps and becomes almost a linear function of time beyond this time, thereby exhibiting diffusive transport as phonon-induced decoherence sets in. For the same exciton-phonon coupling, the MSD appears to be ballistic for a longer time in the case of multilayered material, suggestive of extended coherence lifetime due to the phonon fluctuation synchronization effect discussed above. Thereby, the MSD reaches a higher value in the case of a multilayered material compared to a single-layer material.
@@
Fig. 4b presents MSD in the presence of cavity loss that accounts for the finite lifetime of a photon trapped inside an imperfect optical cavity. Overall, the trends are the same as in Fig. 4a but with an overall lower MSD. These results agree with recent theoretical and experimental works, which demonstrate that MSD is suppressed in the presence of cavity loss [30, 9, 6].
@@
While ballistic transport is suggestive of coherent dynamics, it only provides indirect evidence of decoherence suppression. To provide a quantitative measure and a direct confirmation of decoherence suppression, we compute the purity [45, 46, 47], which is defined as $\mathrm{Tr}[\hat{\rho}_{\mathrm{EP}}^{2}(t)]$ where $\hat{\rho}_{\mathrm{EP}}(t)$ is the reduced density matrix describing the exciton-polariton subsystem. Within our mixed-quantum classical simulations, the exciton-polariton reduced density matrix is computed as $\big{\langle}|\Psi(t)\rangle\langle\Psi(t)|\big{\rangle}_{\mathrm{MFE}}$. In the SI, we provide an efficient approach to computing the purity in our mixed quantum-classical simulations.
@@
Fig. 4c presents the time-dependent purity $\mathrm{Tr}[\hat{\rho}_{\mathrm{EP}}^{2}(t)]$ for initial excitation of $E_{0}=2.6$ eV. In the single-layer scenario (violet solid line), at $t=0$, the purity is one, indicating a pure (coherent) state. At longer times, $t\approx 0.6$ ps, the purity decays toward zero, indicating phonon-induced decoherence of the exciton-polariton subsystem. Fig. 4c, shows that an increase in the number of layers significantly suppresses the decay of the purity, directly depicting an extended coherence lifetime in multilayered materials. Overall, we find that multilayered materials can enhance purity by up to a factor of $\approx 4$ as shown in Fig. 4d, which presents the ratio of purity in a multilayered material relative to a single-layer material.
@@
Fig. 4e-f presents the time-dependent purity in the presence of cavity loss. Note that in the presence of cavity loss, the population in the exciton-photon subspace $\{|1_{k}\rangle,|n,m\rangle\}$ leaks out to the ground state $|\bar{0}\rangle$. As a result, the total reduced density matrix (when tracing over the phonon degree of freedom) can be defined as $\hat{\rho}_{\mathrm{T}}(t)=\hat{\rho}_{\mathrm{g}}(t)+\hat{\rho}_{\mathrm{EP}}(t)=\rho_{\mathrm{gg}}(t)|\bar{0}\rangle\langle\bar{0}|+\sum_{\alpha,\beta}\rho_{\alpha\beta}(t)|\alpha\rangle\langle\beta|$ where $\{|\alpha\rangle,|\beta\rangle\}\in\{|1_{k}\rangle,|n,m\rangle\}$. As a result, the total purity can be written down as
$\mathrm{Tr}[\hat{\rho}_{\mathrm{T}}^{2}(t)]=\mathrm{Tr}[\hat{\rho}_{\mathrm{g}}^{2}(t)]+\mathrm{Tr}[\hat{\rho}_{\mathrm{EP}}^{2}(t)].$ (17)